The measure of the supplement of an angle is 20 degrees more than three times the measure of the original angle. Find the measure of the original angle and the supplement.
Answer by checkley77(12844)

X+(3X+20)=180
X+3X=180-20
4X=160
X=160/4
X=40 DEGREES FOR THE SMALLER ANGLE.
180-40=140 FOR THE LARGER ANGLE.
3*40+20=140
120+20=140